OpenStudy (anonymous):

Solve algebraically... 2/x + 5/3 = 7/x

OpenStudy (anonymous):

put like terms together. \[\frac{2}{x}-\frac{7}{x}=-\frac{5}{3} \] \[\frac{2-7}{x}=-\frac{5}{3}\] and finish on your own.

OpenStudy (anonymous):

x = 3?

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (anonymous):

thanks!